OpenStudy (anonymous):

A line passes through the point (0, –1) and has a slope of -2/3. What is the equation of the line? a y=-2/3x-1 b y=3/2x+1 c y=2/3x+1 d y=-2/3x+1

OpenStudy (shirley27):

are you familiar with the point-slope form of the line formula?

OpenStudy (shirley27):

y-y1 = m * (x-x1) where y1 and x1 are points.

OpenStudy (anonymous):

i think i got it

OpenStudy (lynfran):

the answer y=-2/3x-1

OpenStudy (lynfran):

y-(-1)=-2/3(x-0) y+1=-2/3x y=-2/3x-1